Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared by means of an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

standard quantity standard price or rate standard cost
direct materials 2.5 ounces $20.00 per ounce $50.00
Direct labor 1.4 ounces $12.50 per hour 17.50
67.50

During November, the following activity was recorded by the company relative to production of Fludex:

a. Materials purchased 12,000 ounces at a cost of $225,000.

b. There was no beginning inventory of materials on hand to start the month; at the end of the month, 2,500 ounces of material remained in the warehouse unused.

c. The company employs 35 lab technicians to work on the production of Fludex. During November, each worked an average of 160 hours at an average rate of $12 per hour.

d. During November, 3,750 good units of Fludex were produced.

1) Compute the materials price variance, materials quantity variance, and the total Flexible Budget Materials Variance.

2) Compute the labor rate variance, labor efficiency variance and the total Flexible Budget Labor Variance

3)

a. The materials were purchased from a new supplier who is anxious to enter into a longterm purchase contract. Would you recommend that the company sign the contract? Explain.

b. In the past, the 35 technicians employed in the production of Fludex consisted of 20 senior technicians and 15 assistants. During November, the company experimented with only 15 senior technicians and 20 assistants in order to save costs. Would you recommend that the new labor mix be continued? Explain.

c. What other factors could we infer from the above, or what other questions would you ask to be sure that the correct decision is made?
Answer #1

Solution 1:

Standard quantity of material for actual production = 3750*2.50 = 9375 ounce

Actual quantity of material purchased = 12000 ounce

Actual quantity of material used = 12000 - 2500 = 9500 ounce

Standard price of material = $20 per ounce

Actual price of material = $225,000 / 12000 = $18.75

Material price variance = (SP - AP) * AQ purchased = ($20 - $18.75) * 12000 = $15,000 F

Material quantity variance = (SQ - AQ) * SR = (9375 - 9500) * $20 = $2,500 U

Solution 2:

Standard hours of direct labor = 3750*1.4 = 5250 hours

Standard rate of direct labor = $12.50 per hour

Actual hours of direct labor = 35*160 = 5600 hours

Actual rate of direct labor = $12 per hour

Direct labor rate variance = (SR - AR) * AH = ($12.50 - $12) * 5600 = $2,800 F

Direct labor efficiency variance = (SH - AH) * SR = (5250 - 5600) * $12.50 = $4,375 U

Solution 3a:

As price offered by the new supplier is lesser than standard price of material, therefore company should sign long term purchase contract with the new supplier.

Solution 3b:

Employing more assistant rather senior technician resulted in favorable direct labor rate variance but unfavorable laor efficiency variance. Further unfavorable efficiency variance is higher than favorable rate variance, therefore it is recommended new labor mix should not be continued.

Solution 3c:

Other factors that could be considered are quality of material offered by new supplier, quality of work done by assistant, effect on variable overhead efficiency variance on employing more assistant etc should be considered before making any decision.